# Maximum entropy based testing in network models: ERGMs and constrained optimization

### v1

@@ Appendix A Proofs
$m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ})\xrightarrow{P}\frac{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}{\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}}=\frac{1}{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}.$ (A.9)
@@
$m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ}(p))\xrightarrow{P}\frac{1}{\frac{(u^{*}(p))^{e(H)}}{\lvert Aut(H)\rvert}-\frac{p_{0}^{e(H)}}{\lvert Aut(H)\rvert}}$ (A.10)
@@
$m_{n}^{2}(\tilde{\lambda}_{n}-\lambda^{\circ}(\tilde{p}))\xrightarrow{P}\frac{1}{\frac{(u^{*}(\tilde{p}))^{e(H)}}{\lvert Aut(H)\rvert}-\frac{p_{0}^{e(H)}}{\lvert Aut(H)\rvert}}$ (A.11)
@@
Hence subtracting (A.11) from (A.10) gives the result.
Appendix B Proofs of Theorems 3.8 and 3.10
@@
$f(x)=-\frac{\lambda}{m^{v(H)-2}}H(x)+\log\Bigl(\frac{N(x)}{n}\Bigr).$ (B.1)
@@
In particular we will show that not only the functions are approximated by each other, but the same holds for the first and second derivatives. The term $\log Z_{m}(\bm{\beta})$ is a constant and can be ignored for derivative calculations. Let us denote the right hand side of the above display by $f_{ERGM}$. The function $f(x)$ in (B.1) thus decomposes as
@@
Along with Lemmas C.1, C.2 and C.3, these imply
@@
By Lemma C.2, the same $\epsilon$-net works for the image of $\nabla f$ if $\epsilon$ is not as small as $o_{p}(\frac{1}{m_{n}})$.
@@
$\qquad+\frac{1}{4}\Bigl(\sum_{e}b_{e}^{2}\Bigr)^{1/2}\epsilon+3n_{v}\epsilon+\log|\mathcal{D}(\epsilon)|+S,$ (B.2)
@@
$\log\sum_{x\in\{0,1\}^{\binom{m}{2}}}\exp(f(x))$ $\geq\sup_{x\in[0,1]^{\binom{m}{2}}}\Bigl(f(x)-I(x)\Bigr)-\frac{1}{2}\sum_{e}c_{ee},$ (B.3)
@@
  • $S$ is the smoothness term which, after normalization by $n_{v}$, is $O(m^{-1/2})$ (see Lemma C.4).
@@
After dividing by $m^{2}$, the error terms in the upper bound (B.2) converge to zero. The lower bound error term, $\frac{1}{m^{2}}\sum c_{ee}=\frac{1}{m^{2}}O(n_{v}\cdot B/m)=O(B/m)$, also vanishes. This leads to the following convergence result, stated as a lemma.
Lemma B.1 (Convergence of Partition Function for the Empirical ERGM).
@@
By the uniform weak law of large numbers (Lemma C.1), the ratio inside the logarithm converges to 1 in probability, uniformly over all graphs $x$. Therefore, the error term vanishes as $n\to\infty$:
@@
$=\lim_{m\to\infty}\Biggl[\sup_{x\in[0,1]^{\binom{m}{2}}}\Biggl\{-\frac{\lambda}{m^{v(H)}}H(x)+\sum_{k=1}^{K}\beta_{k}t(T_{k},x)-\frac{I(x)}{m^{2}}\Biggr\}-\frac{1}{m^{2}}Z_{m}(\beta)\Biggr]$
@@
As mentioned in the proof outline we will deal with the numerator and denominator of the RHS of the above display separately. For the numerator we will prove (A.7), while for the denominator we will establish (A.8). Since the statement of the theorem follows from combining (A.7) and (A.8), the rest of the proof will be devoted to proving (A.7) and (A.8). From the proof of Theorem 3.8 we have:
$\frac{1}{m_{n}^{2}}\log(\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\lambda}{m_{n}^{v(H)-2}}h_{n,i}})\xrightarrow{P}\mathfrak{g}(\lambda)$ (B.4)
@@
  • The left hand side of (B.4) is a sequence of convex functions.
  • The right hand side of (B.4), being a pointwise limit of convex functions, is convex.
@@
Thus the convergence outlined in (A.7) is established. Next to analyze convergence in (A.8), we require the limit of
@@
$\frac{1}{m_{n}^{2}}\log\sum_{G}e^{f(G)}+\frac{\lambda p_{0}^{e(H)}}{\lvert Aut(H)\rvert}+\frac{\alpha p_{0}^{2e(H)}}{\lvert Aut(H)\rvert^{2}}$ (B.5)
@@
$\left.+\frac{\alpha}{\lvert Aut(H)\rvert^{2}}\left(\frac{H(x)}{m_{n}^{v(H)}/\lvert Aut(H)\rvert}\right)^{2}\right]$ (B.6)
@@
The energy function in (B.6) now decomposes as
@@
$\frac{1}{m_{n}^{2}}\log\left(\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}}h_{n,i}+\frac{\alpha}{m_{n}^{2v(H)-2}}h_{n,i}^{2}}\right)$ (B.7)
@@
$-\sup_{\tilde{h}\in\tilde{W}}\bigg(\sum_{k=1}^{K}\beta_{k}t(T_{k},h)-\frac{1}{2}I(h)\bigg)$ (B.8)
@@
  • The left hand side of (B.8) is a sequence of convex functions of $\alpha$.
  • The right hand side of (B.8), being a pointwise limit of convex functions, is convex.
  • The left hand side of (B.8) is a sequence of differentiable functions of $\alpha$.
@@
Appendix C Proof of technical lemmas
Lemma C.1 (Uniform Weak Law of Large Numbers).
@@ Appendix C Proof of technical lemmas
Lemma C.2 (Uniform Convergence of the Log-Derivative).
@@
By the uniform weak law Lemma C.1, for every graph $G$ on $m_{n}$ vertices we have
@@
$\mathbb{P}\!\Biggl(\sup_{x\in[0,1]^{\binom{m_{n}}{2}}}\;\sup_{1\leq i<j\leq m_{n}}\;\Bigl|D_{e}(x)-\frac{\partial}{\partial x_{ij}}p_{\beta}(x)|>\frac{\epsilon}{m_{n}}\Biggr)\leq 2^{\binom{m_{n}}{2}}\binom{m_{n}}{2}\,\delta(n).$
@@
Lemma C.3 (Uniform Convergence of the Discrete Second Derivative).
@@
Lemma C.4 (Controlling the smoothness term).
The smoothness term $\mathcal{S}$ in formula (B.2), after normalization by $n_{v}$, is $O(m^{-1/2})$.
@@
$F\leq\sup_{x\in[0,1]^{n_{v}}}(f(x)-I(x))+\text{complexity term}+\text{smoothness term}$ (C.1)
@@
Our term $\mathcal{S}$ in formula (B.2) corresponds to this $S_{\text{NLLD}}$.
@@
$f(x)=-\frac{\lambda}{m^{v(H)-2}}H(x)+\log\left(\frac{N(x)}{n}\right)$ (C.2)
@@
Finally, since the map $\lambda\mapsto\mu(\lambda)=h_{0}e^{\lambda}$, it is easy to observe that $f(\lambda)$ is continuous in $\lambda$.
@@
Defining $Y_{\lambda}:=(Z_{\mu(\lambda)}-h_{0})\,e^{-\lambda(Z_{\mu(\lambda)}-h_{0})}$, we will show that the map $\lambda\longmapsto\mathrm{Var}(Y_{\lambda})$ is continuous. It suffices to prove continuity of $\lambda\mapsto\mathbb{E}[Y_{\lambda}]$ and $\lambda\mapsto\mathbb{E}[Y_{\lambda}^{2}]$.
We have
@@
Appendix D Existence and Uniqueness of roots
@@
To invoke Lemma D.1 and Lemma D.2, we must verify that the support of $Y_{n}$ strictly straddles zero with high probability as $n\to\infty$. That is, we define the event:
@@
    Assume the event $E_{n}$ holds. The support of the empirical variable $Y_{n}$ is $\mathcal{H}_{n}$. By definition of $E_{n}$, we have $\min(\mathcal{H}_{n})<0<\max(\mathcal{H}_{n})$. By Lemma D.2, this condition is necessary and sufficient for the existence of a real root to the equation $\mathbb{E}[Y_{n}e^{-\lambda Y_{n}}]=0$ (and consequently for the scaled equation in the dense regime). Since $\lim_{n\to\infty}\mathbb{P}(E_{n})=1$, the root exists with probability approaching 1.
@@
    Conditioned on $E_{n}$, the values $h_{n,i}$ are not all zero. Therefore, $h_{n,i}^{2}>0$ for at least some indices, implying $f^{\prime}(\lambda)<0$ for all $\lambda\in\mathbb{R}$. The function is strictly decreasing. By Lemma D.1, if a root exists, it must be unique.
Combining the asymptotic probability of $E_{n}$ with the deterministic results of Lemmas D.1 and D.2 conditioned on $E_{n}$: with probability approaching 1 as $n\to\infty$, the Lagrange multiplier $\hat{\lambda}_{n}$ exists and is the unique solution to the estimating equation. ∎

### v2

@@ Appendix C Proofs of the main Theorems
$m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ})\xrightarrow{P}\frac{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}{\left(\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}\right)^{2}}=\frac{1}{\frac{(u^{*})^{e(H)}}{|Aut(H)|}-\frac{p_{0}^{e(H)}}{|Aut(H)|}}.$ (C.9)
@@
$m_{n}^{2}(\hat{\lambda}_{n}-\lambda^{\circ}(p))\xrightarrow{P}\frac{1}{\frac{(u^{*}(p))^{e(H)}}{\lvert Aut(H)\rvert}-\frac{p_{0}^{e(H)}}{\lvert Aut(H)\rvert}}$ (C.10)
@@
$m_{n}^{2}(\tilde{\lambda}_{n}-\lambda^{\circ}(\tilde{p}))\xrightarrow{P}\frac{1}{\frac{(u^{*}(\tilde{p}))^{e(H)}}{\lvert Aut(H)\rvert}-\frac{p_{0}^{e(H)}}{\lvert Aut(H)\rvert}}$ (C.11)
@@
Hence subtracting (C.11) from (C.10) gives the result.
Appendix D Proofs of Theorems 3.8 and 3.10
@@
$f(x)=-\frac{\lambda}{m^{v(H)-2}}H(x)+\log\Bigl(\frac{N(x)}{n}\Bigr).$ (D.1)
@@
In particular we will show that not only the functions are approximated by each other, but the same holds for the first and second derivatives. The term $\log Z_{m}(\bm{\beta})$ is a constant and can be ignored for derivative calculations. Let us denote the right hand side of the above display by $f_{ERGM}$. The function $f(x)$ in (D.1) thus decomposes as
@@
Along with Lemmas E.1, E.2 and E.3, these imply
@@
By Lemma E.2, the same $\epsilon$-net works for the image of $\nabla f$ if $\epsilon$ is not as small as $o_{p}(\frac{1}{m_{n}})$.
@@
$\qquad+\frac{1}{4}\Bigl(\sum_{e}b_{e}^{2}\Bigr)^{1/2}\epsilon+3n_{v}\epsilon+\log|\mathcal{D}(\epsilon)|+S,$ (D.2)
@@
$\log\sum_{x\in\{0,1\}^{\binom{m}{2}}}\exp(f(x))$ $\geq\sup_{x\in[0,1]^{\binom{m}{2}}}\Bigl(f(x)-I(x)\Bigr)-\frac{1}{2}\sum_{e}c_{ee},$ (D.3)
@@
  • $S$ is the smoothness term which, after normalization by $n_{v}$, is $O(m^{-1/2})$ (see Lemma E.4).
@@
After dividing by $m^{2}$, the error terms in the upper bound (D.2) converge to zero. The lower bound error term, $\frac{1}{m^{2}}\sum c_{ee}=\frac{1}{m^{2}}O(n_{v}\cdot B/m)=O(B/m)$, also vanishes. This leads to the following convergence result, stated as a lemma.
Lemma D.1 (Convergence of Partition Function for the Empirical ERGM).
@@
By the uniform weak law of large numbers (Lemma E.1), the ratio inside the logarithm converges to 1 in probability, uniformly over all graphs $x$. Therefore, the error term vanishes as $n\to\infty$:
@@
$=\lim_{m\to\infty}\Biggl[\sup_{x\in[0,1]^{\binom{m}{2}}}\Biggl\{-\frac{\lambda}{m^{v(H)}}H(x)+\sum_{k=1}^{K}\beta_{k}t(T_{k},x)-\frac{I(x)}{m^{2}}\Biggr\}-\frac{1}{m^{2}}\log Z_{m}(\beta)\Biggr]$
@@
As mentioned in the proof outline we will deal with the numerator and denominator of the RHS of the above display separately. For the numerator we will prove (C.7), while for the denominator we will establish (C.8). Since the statement of the theorem follows from combining (C.7) and (C.8), the rest of the proof will be devoted to proving (C.7) and (C.8). From the proof of Theorem 3.8 we have:
$\frac{1}{m_{n}^{2}}\log(\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\lambda}{m_{n}^{v(H)-2}}h_{n,i}})\xrightarrow{P}\mathfrak{g}(\lambda)$ (D.4)
@@
  • The left hand side of (D.4) is a sequence of convex functions.
  • The right hand side of (D.4), being a pointwise limit of convex functions, is convex.
@@
Thus the convergence outlined in (C.7) is established. Next to analyze convergence in (C.8), we require the limit of
@@
$\frac{1}{m_{n}^{2}}\log\sum_{G}e^{f(G)}+\frac{\lambda p_{0}^{e(H)}}{\lvert Aut(H)\rvert}+\frac{\alpha p_{0}^{2e(H)}}{\lvert Aut(H)\rvert^{2}}$ (D.5)
@@
$\left.+\frac{\alpha}{\lvert Aut(H)\rvert^{2}}\left(\frac{H(x)}{m_{n}^{v(H)}/\lvert Aut(H)\rvert}\right)^{2}\right]$ (D.6)
@@
The energy function in (D.6) now decomposes as
@@
$\frac{1}{m_{n}^{2}}\log\left(\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\tilde{\lambda}_{n}}{m_{n}}h_{n,i}+\frac{\alpha}{m_{n}^{2v(H)-2}}h_{n,i}^{2}}\right)$ (D.7)
@@
$-\sup_{\tilde{h}\in\tilde{W}}\bigg(\sum_{k=1}^{K}\beta_{k}t(T_{k},h)-\frac{1}{2}I(h)\bigg)$ (D.8)
@@
  • The left hand side of (D.8) is a sequence of convex functions of $\alpha$.
  • The right hand side of (D.8), being a pointwise limit of convex functions, is convex.
  • The left hand side of (D.8) is a sequence of differentiable functions of $\alpha$.
@@
Appendix E Proof of technical lemmas
Lemma E.1 (Uniform Weak Law of Large Numbers).
@@ Appendix E Proof of technical lemmas
Lemma E.2 (Uniform Convergence of the Log-Derivative).
@@
By the uniform weak law Lemma E.1, for every graph $G$ on $m_{n}$ vertices we have
@@
$\mathbb{P}\!\Biggl(\sup_{x\in[0,1]^{\binom{m_{n}}{2}}}\;\sup_{1\leq i<j\leq m_{n}}\;\Bigl|D_{e}(x)-\frac{\partial}{\partial x_{ij}}\log p_{\beta}(x)|>\frac{\epsilon}{m_{n}}\Biggr)\leq 2^{\binom{m_{n}}{2}}\binom{m_{n}}{2}\,\delta(n).$
@@
Lemma E.3 (Uniform Convergence of the Discrete Second Derivative).
@@
Lemma E.4 (Controlling the smoothness term).
The smoothness term $\mathcal{S}$ in formula (D.2), after normalization by $n_{v}$, is $O(m^{-1/2})$.
@@
$F\leq\sup_{x\in[0,1]^{n_{v}}}(f(x)-I(x))+\text{complexity term}+\text{smoothness term}$ (E.1)
@@
Our term $\mathcal{S}$ in formula (D.2) corresponds to this $S_{\text{NLLD}}$.
@@
$f(x)=-\frac{\lambda}{m^{v(H)-2}}H(x)+\log\left(\frac{N(x)}{n}\right)$ (E.2)
@@
Finally, since the map $\lambda\mapsto\mu(\lambda)=h_{0}e^{\lambda}$ is continuous, it is easy to observe that $f(\lambda)$ is continuous in $\lambda$.
@@
Defining $Y_{\lambda}:=(Z_{\mu(\lambda)}-h_{0})\,e^{-\lambda(Z_{\mu(\lambda)}-h_{0})}$, we will show that the map $\lambda\mapsto\mathrm{Var}(Y_{\lambda})$ is continuous. It suffices to prove continuity of $\lambda\mapsto\mathbb{E}[Y_{\lambda}]$ and $\lambda\mapsto\mathbb{E}[Y_{\lambda}^{2}]$. We have
@@
Appendix F Existence and Uniqueness of roots
@@
To invoke Lemma F.1 and Lemma F.2, we must verify that the support of $Y_{n}$ strictly straddles zero with high probability as $n\to\infty$. That is, we define the event:
@@
    Assume the event $E_{n}$ holds. The support of the empirical variable $Y_{n}$ is $\mathcal{H}_{n}$. By definition of $E_{n}$, we have $\min(\mathcal{H}_{n})<0<\max(\mathcal{H}_{n})$. By Lemma F.2, this condition is necessary and sufficient for the existence of a real root to the equation $\mathbb{E}[Y_{n}e^{-\lambda Y_{n}}]=0$ (and consequently for the scaled equation in the dense regime). Since $\lim_{n\to\infty}\mathbb{P}(E_{n})=1$, the root exists with probability approaching 1.
@@
    Conditioned on $E_{n}$, the values $h_{n,i}$ are not all zero. Therefore, $h_{n,i}^{2}>0$ for at least some indices, implying $f^{\prime}(\lambda)<0$ for all $\lambda\in\mathbb{R}$. The function is strictly decreasing. By Lemma F.1, if a root exists, it must be unique.
Combining the asymptotic probability of $E_{n}$ with the deterministic results of Lemmas F.1 and F.2 conditioned on $E_{n}$: with probability approaching 1 as $n\to\infty$, the Lagrange multiplier $\hat{\lambda}_{n}$ exists and is the unique solution to the estimating equation. ∎